M is a drug recently proven to be effective in treating male pattern baldness. the pharmicist mixes a solution 2%M. He has 50 ml of a 1% solution. He adds 4% solution to it to obtain the desired 2% solution. How much 4% solution does he add?
Answer by Paul(988)

Let the soultion be x
Equation:
4x+1(50)=2(x+50)
4x+50=2x+100
2x=50
x=25
Hence, he should add 25 ml of solution.
Paul.